please help me solve this..i've been working on it for half an hour!
at an ice cream parlor, ice cream cones are $ 1.10 and sundaes are $2.35. one day, the receipts for a total of 172 cones and sundaes were $294.20. how many cones were sold?
thanks so much!!

Found 2 solutions by Paul, whty89:
Answer by Paul(988)

Let ice cream be x
Let cones be y
x+y=172 (total)
solve for y:
y=172-x (subsitutiON)

1.10x+2.35y=294.20
Use subsitution:
1.10x+2.35(172-x)=294.20
1.10x+404.2-2.35x=294.20
-1.25x=-110
x=88

y=172-88
y=84

Hence, there were 88 ice creams and 84 ice cream cones.
Paul.

Answer by whty89(20)

1.) assign variables, x=cones, y=sundaes
2.) set it up as a system of equations, this is how you set it up,
x + y=172
1.1x+2.35y=294.2
3.)solve for a variable by substitution of elimination. I think elimination would be easiest.
4.) multiply the top equation by -1.1 so the x's will cancel, now you have
-1.1x + -1.1y= -189.2, for the top equation
1.1x + 2.35y=294.2, for the bottom equation
5.) you add the two equations together. the x's cancel and you have
1.25y=105, solve for y by dividing 105 by 1.25. y =84
6.) plug 84 back into one of the original equations for y
x + 84 = 172, subtract 84 from both sides x=88
88 cones were sold and 84 sundaes were sold